# Kenyon’s identities for the height function and compactified free field in the dimer model

### v1

@@ Acknowledgment
The author thanks Dmitry Chelkak and Zach Deiman for bringing their very recent research [ChelkakDeiman] to author’s attention before publication and for enlightening discussions of the doubly connected setup. The work was supported by Research Council of Finland grants 333932 and 363549.
@@ 2. Preliminaries on Riemann surfaces
The purpose of this section is to remind some basic facts and constructions from the theory of Riemann surfaces that we will need. We will try to keep the presentation as self-contained as possible, though some of the proofs will have to be omitted. All the facts below are fairly standard and can be found in standard textbooks such as [MumfordTata1, GriffitsHarris, Forster, Fay]; we also address the reader to [basok2023dimers, Appendix] and references therein.
@@ 3.3. Variation of the observable
This follows from the expression for $\operatorname{{\mathbb{E}}}\exp(\pi iQ_{\mathfrak{m}}(\psi-\psi_{0}))$ obtained in [basok2023dimers, Lemma 7.2] and relating it with an appropriate theta constant, and from Riemann theorem on theta divisor, see [basok2023dimers, Theorem 9.3]. ∎
@@
A straightforward calculation shows that the differential $2\pi ir_{\alpha}$ coincides with the differential $W_{\alpha}$ introduced in the proof of [basok2023dimers, Lemma 4.4]. The variation identity now follows from the variation identity given in the proof of [basok2023dimers, Lemma 4.4] and the identity given in [basok2023dimers, Corollary 7.1]. ∎
@@
Fix an arbitrary Hermitian metric on ${\mathcal{F}}_{0}$ and $\Sigma$ and consider the Laplace operator $\Delta_{\beta}={\mathcal{D}}_{\beta}^{\ast}{\mathcal{D}}_{\beta}$. Comparing (3.12) and the classical Quillen’s variation identity [Quillen] one can show that
@@ 3.4. Fermionization of correlations of a compactified free field
Theorem 3.5 allows to give a short proof of determinant expressions for correlations of a compactified free field $h$. In the case of multiply connected domains such expressions appeared in [nicoletti2025temperleyan] (as one can easily see by using the standard formula for ${\mathcal{D}}^{-1}_{\alpha}$ in terms of theta functions [Fay, Chapter VI]).
@@ 5.1. Cylinder with Temperleyan boundaries
To calculate the height change in this case we need to find periods of $\operatorname{\mathrm{Im}}\alpha_{2}$, where $\alpha_{2}$ is such that ${\mathcal{O}}_{\Sigma}(p_{2}-p_{1})\cong{\mathcal{L}}_{\alpha_{2}}$. A standard calculation using Abel–Jacobi map (see e.g. [basok2023dimers, Section 9.4]) implies that
@@ 5.2. Cylinder with black and white boundaries
This setup has been studied in details in [ChelkakDeiman]. Let $\Omega\subset{\mathbb{R}}/{\mathbb{Z}}$ be a doubly connected bounded domain, and let $\Omega_{k}\subset{\mathbb{Z}}^{2}/_{z\sim z+2k}$ be a sequence of discrete doubly connected domains approximating $\Omega$ and such that
@@
Consider now the functions $f^{[\pm,\pm]}$ on $\Omega$ originating from the dimer model on the sequence $\Omega_{k}$. Note that the Kasteleyn operator on each $\Omega_{k}$ can be chosen by taking standard Kasteleyn weights on ${\mathbb{Z}}^{2}$ twisting it with the monodromy 1-1 around the cylinder. Consequently, $f^{[\pm,\pm]}$ will inherit this monodromy 1-1. Summing this with (5.1) (see [ChelkakDeiman] for details) we can find a gauge relating $f^{[\pm,\pm]}$ with the Cauchy kernel in ${\mathcal{F}}_{0}\otimes{\mathcal{L}}$; thus, the height shift is determined by the point in the Jacobian corresponding to $\mathcal{L}$}.
@@ 5.3. Multiply connected domains: Kenyon’s original setup
Let now $\Omega$ be a multiply connected domain with boundary components $B_{0},B_{1},\dots,B_{n},\ n\geq 2,$ where $B_{0}$ denotes the outer component. Let $\Omega_{k}\subset\tfrac{1}{k}{\mathbb{Z}}^{2}$ be discrete domains with Temperleyan boundary conditions approximating $\Omega$. For each $k$ and $i=1,\dots,n$ pick a black vertex $b_{k}$ on the outer boundary of $\Omega_{k}$ and a white vertex $w_{k}^{(i)}$ on the $i$-th boundary component of $\Omega_{k}$. Define $\Omega^{\prime}_{k}=\Omega_{k}\smallsetminus\{b_{k},w_{k}^{(1)},\dots,w_{k}^{(n)}\}$. The dimer model in such $\Omega^{\prime}_{k}$ has been originally considered in the foundational work [KenyonConfInvOfDominoTilings] of Kenyon. Let us show how the result of [nicoletti2025temperleyan] describing the limit of the height function can be re-derived using the language we developed.
@@
Note that from the dimer model perspective removing a white vertex $w_{k}^{(i)}$ is the same as adding an “exposed” black vertex incident to $w_{k}^{(i)}$ and to no other vertices in $\Omega_{k}$. The latter convention was taken in [KenyonConfInvOfDominoTilings].
@@ 5.4. Multiply connected domains: punctures in the bulk
In this example we take the same domains $\Omega_{k}$ as in the previous case but instead of puncturing them at the boundary we remove n1n-1 white vertices $w_{k}^{(1)},\dots,w_{k}^{(n-1)}$ from the bulk of each $\Omega_{k}$. This setup corresponds to the one considered in [basok2023dimers] and [BerestyckiLaslierRayI]. Assume that $w_{k}^{(i)}$ converge to a point $p_{i}\in\Omega$ for each $i=1,\dots,n-1$. The Kasteleyn weights on $\Omega_{k}$ can be chosen such that $f^{[\pm,\pm]}$ are multivalued functions in $\Omega\smallsetminus\{p_{1},\dots,p_{n-1}\}$ with monodromy 1-1 around each inner boundary component and each $p_{1},\dots,p_{n-1}$. Since $\Omega_{k}$ are Temperleyan, $f^{[\pm,\pm]}$ satisfy (5.1).
@@
for some anti-holomorphic $\alpha_{0}$. By construction, ${\mathcal{O}}_{\Sigma}(p_{1}+\ldots+p_{2n-2})$ has a holomorphic section $\omega$ with the divisor $p_{1}+\ldots+p_{2n-2}$. The isomorphism above allows to identify $\omega$ with a multivalued holomorphic differential on $\Sigma$ with the same divisor and monodromy $\chi(\gamma)=\exp(2i\int_{\gamma}\operatorname{\mathrm{Im}}\alpha_{0})$. One can choose $\omega$ and $\alpha_{0}$ such that $\sigma^{\ast}\alpha_{0}=\bar{\alpha}_{0}$ and $\sigma^{\ast}\omega=\bar{\omega}$.
Let us fix simple paths $\gamma_{1},\dots,\gamma_{n-1}$ such that each $\gamma_{i}$ connects $p_{i}$ with $\sigma(p_{i})$ through the outer boundary component of $\Omega$. Given a smooth map $\gamma:[0,1]\to\Sigma\smallsetminus\{p_{1},\dots,p_{2n-2}\}$ define
$q(\gamma)=\frac{1}{2\pi i}\int_{0}^{1}\tfrac{d}{dt}\log\langle\omega(\gamma(t)),\gamma^{\prime}(t)\rangle\,dt-2\pi^{-1}\int_{\gamma([0,1])}\operatorname{\mathrm{Im}}\alpha_{0}+\sum_{j=1}^{n-1}\gamma\cdot\gamma_{j}+1.$
where $\langle\omega(\gamma(t)),\gamma^{\prime}(t)\rangle$ denotes the pairing between the 1-form $\omega(\gamma(t))$ and the tangent vector $\gamma^{\prime}(t)$, and $\gamma\cdot\gamma_{j}$ denotes the algebraic intersection number. Note that if $\omega$ is single valued, then the first term on the right-hand side is nothing but the winding of $\gamma([0,1])$ with respect to $\omega$ as defined above (2.15). In general, it is easy to see that for any loop $\gamma$ the value $q(\gamma)$ is always integer. Moreover, it is possible to show (see [basok2023dimers, Section 9.2]) that $q(\gamma)\mod 2$ descends to a quadratic form on $H_{1}(\Sigma,{\mathbb{Z}}/2{\mathbb{Z}})$, thus there is a spin line bundle $\mathcal{F}$} such that $q=q_{\mathcal{F}}$ (see Section 2.9). Note that $\mathcal{F}$} depends on the choice of $\alpha_{0}$. Note that by definition of $\mathcal{F}$} the multivalued differential $\omega$ induces a multivalued section $\sqrt{\omega}$ of $\mathcal{F}$} over $\Sigma\smallsetminus\{p_{1},\dots,p_{2n-2}\}$ with monodromy along a loop $\gamma$ calculated as
$\chi(\gamma)=\exp(i\int_{\gamma}\operatorname{\mathrm{Im}}\alpha_{0}+\pi i\sum_{j=1}^{n-1}\gamma\cdot\gamma_{j}).$
Now fix a simplicial basis $A_{1},\dots,A_{n},B_{1},\dots,B_{n}$ in $H_{1}(\Sigma,{\mathbb{Z}})$ in the usual way (see Section 3.1, note that $B_{i}$-s are inner boundary components of $\Omega$), and consider the spin line bundle ${\mathcal{F}}_{0}$ with zero characteristics. Comparing quadratic forms of $\mathcal{F}$} and ${\mathcal{F}}_{0}$ we conclude that every holomorphic section of $\mathcal{F}$} can be identified with a multivaued holomorphic section of ${\mathcal{F}}_{0}$ over $\Sigma$ with monodromy 1-1 along each $B_{i}$ and monodromy $s_{i}=\pm 1$ along each $A_{i}$. Adjusting $\alpha_{0}$ we can ensure that $s_{1}=\ldots=s_{n}=1$. It follows that $\sqrt{\omega}$ can be identified with a multivalued section of ${\mathcal{F}}_{0}$ with monodromy
$\chi_{0}(\gamma)=\exp(i\int_{\gamma}\operatorname{\mathrm{Im}}\alpha_{0}+\pi i\sum_{j=1}^{n-1}\gamma\cdot\gamma_{j}+\pi i\sum_{j=1}^{n}\gamma\cdot A_{j}).$
Assume now that the line bundle ${\mathcal{F}}_{0}\otimes{\mathcal{L}}_{\tfrac{\alpha_{0}}{2}}$ has a Cauchy kernel ${\mathcal{D}}^{-1}$. Identified this kernel with a multivalued section of ${\mathcal{F}}_{0}\boxtimes{\mathcal{F}}_{0}$ as usual and define
@@
It is easy to see that $f^{[\pm,\pm]}$ defined in this way satisfy all the required properties. We conclude that the shift of the compactified free field $h$ is given by $\pi^{-1}\operatorname{\mathrm{Im}}\alpha_{0}$. This agrees with the results of [basok2023dimers].
@@
  • This differential has a geometric meaning: indeed, consider the Riemannian metric given by $|\omega|^{2}$, then this is the unique (up to a multiplicative constant) locally flat metric on $\Sigma$ with conical singularities of angles $4\pi$ at $p_{1},\dots,p_{2n-2}$. The character $\chi(\gamma)=\exp(2i\int_{\gamma}\operatorname{\mathrm{Im}}\alpha_{0})$ is nothing but the holonomy of this metric. This appearance of the holonomy is very consonant with the approach of [BerestyckiLaslierRayI] where the dimer height function is related with the winding of branches of random trees, see [basok2023dimers] for more details.
  • It is worth noting that the construction above fix the square root $\frac{1}{2}\alpha_{0}]\in\operatorname{\mathrm{Jac}}(\Sigma)$ of $\alpha_{0}]\in\operatorname{\mathrm{Jac}}(\Sigma_{0})$, that is, it fixes $2\pi^{-1}\operatorname{\mathrm{Im}}\alpha_{0}$ up to an integer differential with even periods. Note that this may force $\alpha_{0}$ to be non-zero even when the holonomy of the metric above it trivial. This, for example, happens in the case of a Temperleyan cylinder discussed in the first example: in this case $2\pi^{-1}\operatorname{\mathrm{Im}}\alpha_{0}$ is required to have an odd $A$-period which results in the half-integer shift of the height change.

### v2

@@ Acknowledgment
The author thanks Dmitry Chelkak and Zach Deiman for bringing their very recent research [9] to author’s attention before publication and for enlightening discussions of the doubly connected setup. The work was supported by Research Council of Finland grants 333932 and 363549.
@@ 2. Preliminaries on Riemann surfaces
The purpose of this section is to remind some basic facts and constructions from the theory of Riemann surfaces that we will need. We will try to keep the presentation as self-contained as possible, though some of the proofs will have to be omitted. All the facts below are fairly standard and can be found in standard textbooks such as [24, 17, 15, 14]; we also address the reader to [2, Appendix] and references therein.
@@ 3.3. Variation of the observable
This follows from the expression for $\operatorname{{\mathbb{E}}}\exp(\pi iQ_{\mathfrak{m}}(\psi-\psi_{0}))$ obtained in [2, Lemma 7.2] and relating it with an appropriate theta constant, and from Riemann theorem on theta divisor, see [2, Theorem 9.3]. ∎
@@
A straightforward calculation shows that the differential $2\pi ir_{\alpha}$ coincides with the differential $W_{\alpha}$ introduced in the proof of [2, Lemma 4.4]. The variation identity now follows from the variation identity given in the proof of [2, Lemma 4.4] and the identity given in [2, Corollary 7.1]. ∎
@@
Fix an arbitrary Hermitian metric on ${\mathcal{F}}_{0}$ and $\Sigma$ and consider the Laplace operator $\Delta_{\beta}={\mathcal{D}}_{\beta}^{\ast}{\mathcal{D}}_{\beta}$. Comparing (3.12) and the classical Quillen’s variation identity [26] one can show that
@@ 3.4. Fermionization of correlations of a compactified free field
Theorem 3.5 allows to give a short proof of determinant expressions for correlations of a compactified free field $h$. In the case of multiply connected domains such expressions appeared in [25] (as one can easily see by using the standard formula for ${\mathcal{D}}^{-1}_{\alpha}$ in terms of theta functions [14, Chapter VI]).
@@ 5.1. Cylinder with Temperleyan boundaries
To calculate the height change in this case we need to find periods of $\operatorname{\mathrm{Im}}\alpha_{2}$, where $\alpha_{2}$ is such that ${\mathcal{O}}_{\Sigma}(p_{2}-p_{1})\cong{\mathcal{L}}_{\alpha_{2}}$. A standard calculation using Abel–Jacobi map (see e.g. [2, Section 9.4]) implies that
@@ 5.2. Cylinder with black and white boundaries
This setup has been studied in details in [9]. Let $\Omega\subset{\mathbb{R}}/{\mathbb{Z}}$ be a doubly connected bounded domain, and let $\Omega_{k}\subset{\mathbb{Z}}^{2}/_{z\sim z+2k}$ be a sequence of discrete doubly connected domains approximating $\Omega$ and such that
@@
Consider now the functions $f^{[\pm,\pm]}$ on $\Omega$ originating from the dimer model on the sequence $\Omega_{k}$. Note that the Kasteleyn operator on each $\Omega_{k}$ can be chosen by taking standard Kasteleyn weights on ${\mathbb{Z}}^{2}$ twisting it with the monodromy 1-1 around the cylinder. Consequently, $f^{[\pm,\pm]}$ will inherit this monodromy 1-1. Summing this with (5.1) (see [9] for details) we can find a gauge relating $f^{[\pm,\pm]}$ with the Cauchy kernel in ${\mathcal{F}}_{0}\otimes{\mathcal{L}}$; thus, the height shift is determined by the point in the Jacobian corresponding to $\mathcal{L}$}.
@@ 5.3. Multiply connected domains: Kenyon’s original setup
Let now $\Omega$ be a multiply connected domain with boundary components $B_{0},B_{1},\dots,B_{n},\ n\geq 2,$ where $B_{0}$ denotes the outer component. Let $\Omega_{k}\subset\tfrac{1}{k}{\mathbb{Z}}^{2}$ be discrete domains with Temperleyan boundary conditions approximating $\Omega$. For each $k$ and $i=1,\dots,n$ pick a black vertex $b_{k}$ on the outer boundary of $\Omega_{k}$ and a white vertex $w_{k}^{(i)}$ on the $i$-th boundary component of $\Omega_{k}$. Define $\Omega^{\prime}_{k}=\Omega_{k}\smallsetminus\{b_{k},w_{k}^{(1)},\dots,w_{k}^{(n)}\}$. The dimer model in such $\Omega^{\prime}_{k}$ has been originally considered in the foundational work [22] of Kenyon. Let us show how the result of [25] describing the limit of the height function can be re-derived using the language we developed.
@@
Note that from the dimer model perspective removing a white vertex $w_{k}^{(i)}$ is the same as adding an “exposed” black vertex incident to $w_{k}^{(i)}$ and to no other vertices in $\Omega_{k}$. The latter convention was taken in [22].
@@ 5.4. Multiply connected domains: punctures in the bulk
In this example we take the same domains $\Omega_{k}$ as in the previous case but instead of puncturing them at the boundary we remove n1n-1 white vertices $w_{k}^{(1)},\dots,w_{k}^{(n-1)}$ from the bulk of each $\Omega_{k}$. This setup corresponds to the one considered in [2] and [3]. Assume that $w_{k}^{(i)}$ converge to a point $p_{i}\in\Omega$ for each $i=1,\dots,n-1$. The Kasteleyn weights on $\Omega_{k}$ can be chosen such that $f^{[\pm,\pm]}$ are multivalued functions in $\Omega\smallsetminus\{p_{1},\dots,p_{n-1}\}$ with monodromy 1-1 around each inner boundary component and each $p_{1},\dots,p_{n-1}$. Since $\Omega_{k}$ are Temperleyan, $f^{[\pm,\pm]}$ satisfy (5.1).
@@
for some anti-holomorphic $\alpha_{0}$. By construction, ${\mathcal{O}}_{\Sigma}(p_{1}+\ldots+p_{2n-2})$ has a holomorphic section $\omega$ with the divisor $p_{1}+\ldots+p_{2n-2}$. The isomorphism above allows to identify $\omega$ with a multivalued holomorphic differential on $\Sigma$ with the same divisor and monodromy $\chi(\gamma)=\exp(-2i\int_{\gamma}\operatorname{\mathrm{Im}}\alpha_{0})$. One can choose $\omega$ and $\alpha_{0}$ such that $\sigma^{\ast}\alpha_{0}=\bar{\alpha}_{0}$ and $\sigma^{\ast}\omega=\bar{\omega}$.
Let us fix simple paths $\gamma_{1},\dots,\gamma_{n-1}$ such that each $i$ the path $\gamma_{i}$ connects $p_{i}$ with $\sigma(p_{i})$ and $\sigma(\gamma_{i})=\gamma_{i}$. Fix a simplicial basis $A_{1},\dots,A_{n},B_{1},\dots,B_{n}$ in $H_{1}(\Sigma,{\mathbb{Z}})$ in the usual way (see Section 3.1, note that $B_{i}$-s are inner boundary components of $\Omega$). Let ${\mathcal{F}}_{0}$ be the spin line bundle with zero characteristics. By the definition of the quadratic form associated with a spin line bundle (see Section 2.9) we can identify $\sqrt{\omega}$ with a multivalued section of ${\mathcal{F}}_{0}$ over $\Sigma\smallsetminus\{p_{1},\dots,p_{2n-2}\}$ whose monodromy around a loop $\gamma$ is given by
(5.4) $\chi(\gamma)=-\exp\left(\pi i(\operatorname{\mathrm{wind}}(\gamma,\omega)-q_{0}(\gamma))\right)=\pm\exp\left(-i\int_{\gamma}\operatorname{\mathrm{Im}}\alpha_{0}+\pi i\sum_{j=1}^{n-1}\gamma\cdot\gamma_{j}\right)$
where $\operatorname{\mathrm{wind}}$ is defined as in Section 2.9 and the choice of $\pm$ in front of the exponent depend on the choice of $\alpha_{0}$. The symmetries of $\omega$ and $\alpha_{0}$ with respect to the involution $\sigma$ force the sign to be 1-1 if $\gamma=B_{i}$ for some $i$. Adjusting $\alpha_{0}$ we can also achieve the sign to be +1+1 when going around $A_{i}$ for each $i$.
Assume now that the line bundle ${\mathcal{F}}_{0}\otimes{\mathcal{L}}_{\tfrac{-\alpha_{0}}{2}}$ has a Cauchy kernel ${\mathcal{D}}^{-1}$. Identify this kernel with a multivalued section of ${\mathcal{F}}_{0}\boxtimes{\mathcal{F}}_{0}$ as usual and define
@@
It is easy to see that $f^{[\pm,\pm]}$ defined in this way satisfy all the required properties. We conclude that the shift of the compactified free field $h$ is given by $\pi^{-1}\operatorname{\mathrm{Im}}\alpha_{0}$. This agrees with the results of [2].
@@
  • This differential has a geometric meaning: indeed, consider the Riemannian metric given by $|\omega|^{2}$, then this is the unique (up to a multiplicative constant) locally flat metric on $\Sigma$ with conical singularities of angles $4\pi$ at $p_{1},\dots,p_{2n-2}$. The character $\chi(\gamma)=\exp(2i\int_{\gamma}\operatorname{\mathrm{Im}}\alpha_{0})$ is nothing but the holonomy of this metric. This appearance of the holonomy is very consonant with the approach of [3] where the dimer height function is related with the winding of branches of random trees, see [2] for more details.
  • It is worth noting that the construction above fixes the square root $\frac{1}{2}\alpha_{0}]\in\operatorname{\mathrm{Jac}}(\Sigma)$ of $\alpha_{0}]\in\operatorname{\mathrm{Jac}}(\Sigma_{0})$, that is, it fixes $\pi^{-1}\operatorname{\mathrm{Im}}\alpha_{0}$ up to an integer differential with even periods. Note that this may force $\alpha_{0}$ to be non-zero even when the holonomy of the metric above it trivial. This, for example, happens in the case of a Temperleyan cylinder discussed in the first example: in this case $\pi^{-1}\operatorname{\mathrm{Im}}\alpha_{0}$ is required to have an odd $A$-period which results in the half-integer shift of the height change.